A runner starts from rest and accelerates uniformly to a speed of 8.0 meters per second in 4.0 seconds. What is the magnitude of the acceleration of the runner?
1
Expert's answer
2020-09-14T10:19:08-0400

solution;

given data

initial velocity(u)=0m/s

final velocity(v)=8m/s

time(t)=4sec


according to first equation of motion

"v=u+at"

acceleration

"a=\frac{v-u}{t}"

putting the value


"a=\frac{8-0}{4}=2m/s^2"


therefore acceleration of runner is "2m\/s^2" .